A mother is now 3 times added as her daughter. In 4 years the mother will be 4 times s old as her daughter was 2years ago. How old they now?
Answer by lwsshak3(11628)

A mother is now 3 times added as her daughter. In 4 years the mother will be 4 times s old as her daughter was 2years ago. How old they now?
***
let x=daughters present age
3x=mother's present age
..
2yrs ago:
x-2=daughter's age
..
In 4 yrs:
3x+4=mother's age
3x+4=4(x-2)
3x+4=4x-8
x=12
3x=36
daughters present age=12
mother's present age=36